# Experimental Exploration: Investigating Cooperative Interaction Behavior Between Humans and Large Language Model Agents

### v1

@@ 5.1. Behind the Masks: Cooperation Behavior
Male P vs. Purported Human - Male P vs. Purported Rule-Based Agents .1600 4.484 < .001 ***
Male P vs. Purported Human - Male P vs. Purported LLM Agents .0733 2.055 .0439 *
Male P vs. Purported Rule-Based AI Agents - Male P vs. Purported LLM Agents -.0867 -2.429 .0179 *
Male P vs. Purported Rule-Based Agents - Female P vs. Purported Rule-Based AI Agents -.1027 -2.100 .0395 *
Female P vs. Purported Rule-Based AI Agents - Female P vs. Purported LLM Agents .0840 2.354 .0216 *
In the table above, asterisks (*) denote the statistical significance level of the results presented. Specifically: *: p < 0.05, indicating a significant difference between groups or conditions at the 5% level; **: p < 0.01, indicating a highly significant difference between groups or conditions at the 1% level; ***: p < 0.001, indicating an extremely significant difference between groups or conditions at the 0.1% level. These significance levels highlight the strength of the evidence against the null hypothesis, with more asterisks indicating stronger evidence of a true effect.
There are six combinations of interaction effects, Figure 4. Additionally, Table 6 also listed the combinations with significant differences, which revealed that female participants had a higher cooperation rate ($p$ = .0216) cooperate with the purported rule-based AI agent ($Md$ = .52, $SD$ = .1341) and the LLM agent ($Md$ = .44, $SD$ = .1479). The results also revealed significant differences in the cooperation rate of male participants when they faced different opponents:
@@
  • Male participants had a higher cooperation rate ($p$ = .0439) when interacting with the purported human than a LLM agent ($Md$ = .54, $SD$ = .1244).
  • Male participants had a higher cooperation rate ($p$ = .0179) when interacting with an LLM agent than a purported rule-based AI agent.
  • Female participants had a higher cooperation rate ($p$ = .0395) than male participants when interacting with a purported rule-based AI agent.
Contrary to the results observed among female participants, male participants preferred cooperating with the LLM agents over rule-based AI agents. The underlying reasons for this preference remain unclear; nevertheless, this constitutes a novel and noteworthy finding that warrants further investigation. Additionally, we found that male participants exhibited variations in their cooperative behavior when confronted with opponents of three different purported characteristics, and these differences were statistically significant. Contrary to the results observed among female participants, male participants preferred cooperating with the LLM agents over rule-based AI agents. The underlying reasons for this preference remain unclear; nevertheless, this constitutes a novel and noteworthy finding that warrants further investigation. Additionally, we found that male participants exhibited variations in their cooperative behavior when confronted with opponents of three different purported characteristics, and these differences were statistically significant.
5.2. Think Twice: Longer Deliberations with Humans
Figure 5. Completion Time by Gender and Purported Opponents’ Characteristic
The gender of the participant ($p$ < .001), the purported characteristic of the opponents ($p$ < .001), and the interactions between these factors ($p$ = .0189) all significantly influenced the time taken to complete the game from the LME regression results. We illustrate the relationships among these variables using heatmap diagrams, which depict the time consumed by different participants when completing the same game against opponents of various purported characteristics. Please refer to Figure 5 for a detailed visualization. The data has been sorted in ascending order of completion time for clarity. Our observations revealed a consistent trend among all participants: the longest completion times were recorded when playing against opponents purported to be human, followed by LLM agents, with the shortest completion times occurring against rule-based AI agents. The primary reason for this phenomenon is the varying amount of deliberation time participants engaged in before making their choices. Another notable finding is that male participants took significantly longer to complete the game (all three purported opponents’ characteristics) than female participants, suggesting they spent more time analyzing their strategies.
5.3. Human Generosity: Favorable Behavior in Human-AI Dynamics
Figure 6. Amicability by Gender and Purported Opponents’ Characteristic
Table 4 demonstrated a significant interaction effect between participants’ gender and the purported characteristic on participants’ favourable behavior. This result indicates a selective tendency in favourable behavior towards opponents with varying characteristics under competitive scenarios. Therefore, we computed estimated marginal means (EMMs) for each combination of participants’ gender and opponents’ purported characteristics.Figure 6 demonstrated all significant difference between the two groups. Female participants displayed significantly more favourable behavior($p$ = .0064) towards the purported rule-based AI agent ($Md$ = .4286, $SD$ = .1743) compared to the LLM agent ($Md$ = .35, SD = .1619). Male participants also displayed significantly more favourable behavior ($p$ = .0305) towards the purported human agent ($Md$ = .4167, $SD$ = .083) than a purported rule-based AI Agent. Additionally, female participants were significantly more inclined ($p$ = .0055) to show friendliness towards the purported rule-based AI agent compared to male participants ($Md$ = .3333, $SD$ = .1419).
5.4. Let bygones be bygones: Repairing Relationships
Figure 7. Repair Effort by Purported Opponents’ Characteristic
Table 4 indicates that the purported characteristic of the opponents significantly affects participants’ acceptance of the opponents’ favourable behavior and the subsequent rebuilding of cooperation. Therefore, we constructed a Linear Mixed-Effects (LME) model using only the purported characteristic as the variable, and the results are presented. Unlike amicability, human acceptance is directly proportional to the amount of repair effort. Higher levels of repair effort signify a greater willingness to accept favourable behavior and an increased likelihood of re-establishing bilateral cooperation. Please refer to Figure 7. We found that opponents purported to be human ($Md$ = .4226, $SD$ = .0976) were the most readily accepted by participants for their overtures and in establishing cooperative relationships. This phenomenon was particularly pronounced ($p$ = .0078) when interacting with LLM agents ($Md$ = .35, $SD$ = .1261) and marginally significant ($p$ = .0557) when dealing with rule-based AI agents ($Md$ = .3875, $SD$ = .1157).
5.5. Verification of Hypotheses
In this section, we summarize the key findings of our study and evaluate their implications for the proposed hypotheses. By analyzing the results, we assess whether the observed data support or refute our initial hypotheses.
Table 7. Findings Summary and Hypothesis Verification
Findings H1 H2 H3
Gender significantly affects participants’ game completion time. Support
Purported characteristic significantly affects participants’ cooperation rate. Support
Purported characteristic significantly affects participants’ completion time. Support
Purported characteristic significantly affects participants’ acceptance of favourable behavior. Support
Interaction effect significantly affects participants’ cooperation rate. Support
Interaction effect significantly affects participants’ cooperation rate. Support
Interaction effect significantly affects participants’ completion time. Support
Interaction effect significantly affects participants’ favourable tendency. Support
Male participants had higher cooperation rate with the purported human compared with the rule-based AI agents. Support Support
Male participants had higher cooperation rate with the purported human compared with the LLM agents. Support Support
Male participants had higher cooperation rate with the LLM agents compared with the rule-based AI agents. Support Support
Female participants had higher cooperation rate with the rule-based AI agents than male participants. Support Support
Female participants had higher cooperation rate with the rule-based AI agents compared with the LLM agents. Support Support
Male participants spent more time interacting with opponents of the same purported characteristics than female participants. Support Support
Participants engaged longest with purported human, followed by LLM agents, and shortest with rule-based AI agents. Support Support
Female participants exhibited more favorable behavior toward rule-based AI agents than toward LLM agents. Support Support
Male participants exhibited more favorable behavior toward LLM agents than toward rule-based AI agents. Support Support
Female participants exhibited more favorable behavior toward rule-based AI agents than male participants. Support Support
Participants had higher level of acceptance of repair effort toward purported human than LLM agents. Support Support
Participants had higher level of acceptance of repair effort toward purported human than rule-based AI agents. Support Support
We analyzed our hypotheses from the perspectives of four metrics. The resulting conclusions, presented in Table 7, support our proposed hypotheses. Our final findings indicate that the purported characteristics of AI agents significantly influence human cooperative behavior with them in competitive scenarios across multiple dimensions. Additionally, the gender of human participants affects cooperative behavior in certain contexts, although this effect is not statistically significant. Importantly, the interaction between gender and AI agents significantly impacts human cooperative behavior, exhibiting distinct behavioral pattern characteristics. These findings validate our hypotheses and provide empirical evidence.
5.6. Purported Human Versus Actual Human
Figure 8. Difference Between Purported Human and Human Opponents
@@ 5.7. Questionnaires
Figure 9. Questionnaire Statistical Result
The survey questionnaire encompasses questions regarding attitudes towards experimental content (Q1,Q4,Q5) and past experiences (Q2, Q3) with the use of AI tools and applications. Please refer to the Figure 9 for the specific questions. Regarding Q1, collected data indicates that human players frequently employ punitive and rewarding measures when engaging with the AI Agent in the experiment. This behavior parallels interaction patterns typical of human-to-human engagements, wherein emotional responses precipitate the administration of rewards and punishments (Wu et al., 2022; Molleman et al., 2019). Compared to rule-based AI agents, 86.7% of human players more frequently apply rewards and punishments to the LLM Agent. Q4 inquired about the mindset of human players during the experimental process. Despite our initial clarification that participants should focus solely on maximizing their scores, without considering their opponents’ rewards, they still expected to achieve higher scores than their opponents. This observation suggests that if AI becomes an integral part of society rather than merely a tool for humans, users’ psychological motivations (such as a sense of achievement, self-esteem, and other intrinsic factors) will significantly influence their behavioral tendencies. To some extent, this corroborates the perspective of social comparison theory (Wood, 1996; Suls et al., 2002), highlighting an important consideration for human-AI interaction in future societal contexts. This also indicates that more complex and seemingly autonomous artificial intelligence can enhance human intrinsic motivation to surpass it, thereby gaining a greater sense of achievement. These results further validate why participants spent more time against the LLM agent than the purported rule-based AI agent during the experiment.
In the statistical analysis of Q5, we observed that participants expressed a higher degree of moral condemnation towards the LLM agent than the rule-based AI agent when confronted with betrayal by the opponents. We asked participants who chose the LLM agent about their reasons and summarized several potential underlying factors as follows:
@@
For Q2, more respondents believed that the results produced by rule-based AI agents were more reliable. Regarding Q3, we posited that this preference is related to the participants’ usual work environment. For instance, artists tend to prefer LLMs for providing novel creative ideas, whereas engineers favor rule-based AI agents for delivering precise data or analytical outcomes. The evidence presented in this section substantiates the validation of H2.
6. Semi-structured Interview
The qualitative analysis of participants’ behaviors through semi-structured interviews is grounded in two key reasons. First, while objective game data provide valuable insights, semi-structured interviews allow for a deeper exploration of the underlying reasons behind participants’ behaviors, rather than merely capturing isolated snapshots. Second, existing research has not identified the factors leading to specific experimental outcomes. As such, the inductive interview results offer a foundation for uncovering potential causes and advancing studies on human-AI interaction behaviors in competitive contexts.
To achieve these objectives, the semi-structured interview protocol focuses on participants’ strategic decisions when interacting with LLM agents of varying purported characteristics and their responses to non-cooperative behaviors. The researchers posed two key questions: (1) What was your strategy against the three opponents, and why? (2) How did you feel and adjust your strategy when facing non-cooperative behaviors (defection)? Participants reviewed their recorded game data to aid recollection and contextualize their strategic choices. All interviews were audio-recorded in English, and lasted 15 minutes per participant following each of the first three games. This approach facilitated a comprehensive exploration of participants’ cooperative strategies and their emotional and strategic responses to non-cooperative actions.
Interviews were transcribed by the first author within 48 hours of each session. During the transcription process, nonessential fillers and disfluencies were removed. The transcribed data were manually coded using spreadsheets, highlighting tools, and affinity mapping, employing a thematic coding process. To minimize bias, two authors collaboratively conducted two rounds of initial coding. The coding process followed an iterative approach, involving two rounds of coding to merge, split, or refine preliminary codes. After finalizing the codes, a separate meeting was held among three authors to consolidate these basic codes into themes that directly addressed the research questions. Ultimately, two major themes were identified, encompassing three distinct sub-themes to support the study’s objectives.
6.1. Cooperative Behavior Strategies
@@ 6.1.2. Rationalism in Cooperative Behavior Toward Rule-Based AI Agents
When confronted with a purported rule-based AI agent, participants indicated they would begin to betray earlier to probe the opponent’s punishment strategy, thereby adapting their behavior. P4 stated, "It cannot actually obtain money; it is merely following the strategy programmed by its developers." P14 stated, "I believe I can identify its pattern, and I will formulate my strategy based on its inherent logic." Participants believed that the rule-based AI agent was predictable and that they could discern the underlying strategy through experimentation. The current perception of rule-based AI in the human mind is that it remains predictable and distinguishable. Consequently, when cooperating with such systems, participants tend to use early variations in strategy to probe the behavior patterns of rule-based AI. This enables them to formulate strategies based on the observed responses, reflecting a behaviorist approach.
@@ 6.1.3. Gradual and Persuasive Guidance: Influencing LLM with My Behavior
6.2. Emotion Following Non-cooperative and Its Impact on Subsequent Behavior
6.2.1. Empathy for Human
Participants tend to adopt an understanding and forgiving attitude towards the purported human’s betrayal, exhibiting greater empathy in such scenarios. Moreover, they are inclined to spend more time analyzing the reasons behind the opponent’s choice to betray and considering the subsequent strategies that may follow. For instance, P18 remarked, “I can understand his/her decision making, human nature is all complicated, a strategy to get more bang for your buck I guess.” Similarly, P16 stated, “I would forgive him/her if it wasn’t a recurrent betrayal, and I would be willing to put him/her on probation to work together again.” The findings indicate that participants exhibit greater empathy and understanding towards betrayals by purported humans, recognizing the complexity of human characteristics.
6.2.2. The Betrayal is Programmers’ Decisions
Given the participants’ indifference towards betrayals by rule-based AI agents, it is evident that they attribute these actions to the deterministic characteristic of code logic rather than any form of intentionality. This perception minimizes their emotional response to such betrayals. As P17 aptly remarked, “I don’t feel anything; it’s just the fixed choices of code.” P21 stated, “What algorithm is this giving it decisions? Is it an expert algorithm?.”The content demonstrates that participants exhibit minimal emotional reactions to betrayals by rule-based AI agents, understanding these actions as outcomes of predefined algorithms authored by humans. This underscores a clear distinction in their perceptions and emotional engagements when comparing betrayals by AI agents to those by humans, highlighting the importance of perceived intentionality and human agency in eliciting empathetic and complex emotional responses.
6.2.3. The Betrayal of LLMs Leaves Me Apprehensive

### v2

@@ 4.1.2. Decision Latency: Purported Human > Purported LLM Agents > Purported Rule-based AI Agents (RQ1)
Purported Human vs. Purported Rule-Based Agents 80.1000 7.051 <.001 ***
Purported Human vs. Purported LLM Agents 55.5333 4.888 <.001 ***
Purported LLM Agents vs. Purported Rule-Based Agents 24.5667 2.163 .0347 *
Purported Human: Emmean =454.167, 95%CI = [428.090, 480.2431]; Purported Rule-based Agents: Emmean =374.067, 95%CI = [347.990, 400.143]; Purported LLM Agents: Emmean =398.633, 95% = [372.557, 424.710]
4.1.3. Trust Restoration Tolerance: Purported Human > Purported LLM Agents (RQ1)
We constructed a Linear Mixed-Effects model using only the declared identity as the variable, and the results are presented in Table 6. It reveal that participants exhibited a significantly higher trust restoration tolerance when interacting with agents identified as purported humans, compared to LLM agents ($\beta$= .0770, t=2.750, $p$ = .0079).
Table 6. Declared Identity Contrast on Trust Restoration Tolerance
Contrast Estimate t Ratio Pr(>F) Sig
Purported Human vs. Purported Rule-Based Agents 0.0546 1.950 .0560 -
Purported Human vs. Purported LLM Agents 0.0770 2.750 .0079 *
Purported LLM Agents vs. Purported Rule-Based Agents -0.0224 -0.799 .4274 -
Purported Human: Emmean = .4357, 95%CI = [ .3944, .4770]; Purported Rule-based Agents: Emmean = .3810, 95%CI = [ .3400, .4224]; Purported LLM Agents: Emmean = .3587, 95% = [ .3174, .4000]
4.1.4. Decision Latency: Male > Female (RQ2)
We conducted the linear mixed-effects model to examine the influence of the participant gender on the Decision Latency. Base on the result in Table 7, the analysis revealed a statistically significant difference in decision latency between male and female participants, with females showing shorter decision-making times compared to males ($\beta$=-67.0222, t=-3.536, $p$ = .0014).
Table 7. Gender Contrast on Decision Latency
Contrast Estimate t Ratio Pr(>F) Sig
Female vs. Male -67.0222 -3.536 .0014 **
Female: Emmean =375.444, 95%CI = [347.992, 402.897]; Male: Emmean =442.467, 95%CI = [415.014, 469.919]
4.2. The Interaction Effect of AI Agent’s Declared Identity and Gender
We identified a significant interaction effect between gender and the AI agent’s declared identity, which has a notable impact on certain metrics. To provide a more comprehensive understanding of these disparities, we conducted additional data analyses specifically targeting these metrics. The linear mixed-effects model was also employed to investigate the relationship between the dependent variable and the six group combinations, which were derived from the post-hoc verification of results provided in the Section  4.1.
In these models, we mainly focused on comparing (1) the differences in cooperative behaviors and attitudes towards three types of declared AI identities within a certain gender level, and (2) the differences in cooperative behaviors and attitudes from two types of participants within a certain declared AI identity level.
4.2.1. Cooperation Rate
The results are visualized in Figure 3, which provides a detailed representation of the observed trends and group differences of cooperation rate.
@@ 4.2.1. Cooperation Rate
With reference to Section  4.1.1, we found that the significantly higher cooperation rate in the purported human condition compared to the purported LLM agents and purported rule-based AI agents conditions is primarily attributed to the pronounced difference observed among male participants. Specifically, males exhibited a substantially higher cooperation rate toward purported human, which served as the main driver of this disparity.
Figure 3. Cooperation Rate by AI Agent’s Declared Identity and Gender
4.2.2. Decision Latency
The results are visualized in Figure 4, which provides a detailed representation of the observed trends and group differences of decision latency.
@@ 4.2.2. Decision Latency
With reference to Section  4.1.2, we found that the significantly longer decision latency in the purported human condition compared to the purported LLM agents and purported rule-based AI agents conditions is attributed to the pronounced difference observed among all participants. Specifically, we also found that when interacting with AI agents of three different declared identities, male participants exhibited longer decision latencies compared to female participants.
Figure 4. Decision Latency by Declared Identity of AI Agent and Gender
4.2.3. Unsolicited Cooperation
The results are visualized in Figure 5, which provides a detailed representation of the observed trends and group differences of unsolicited cooperation.
[…]
Figure 5. Unsolicited Cooperation Acts by Declared Identity and Gender
@@ 4.3. Questionnaires
Regarding Q1: The results show that human participants frequently apply rewards and punishments when interacting with the AI agent, resembling patterns seen in human-to-human interactions. Specifically, 86.7% of participants were more likely to use such measures with the LLM Agent compared to rule-based AI agents. For Q2, 60.0% participants believed that the results produced by rule-based AI agents were more reliable. Regarding Q3, the responses provided by the participants were evenly divided. We posited that this preference is related to the participants’ usual work environment. For instance, artists tend to prefer LLMs for providing novel creative ideas, whereas engineers favor rule-based AI agents for delivering precise data or analytical outcomes. For Q4, 63.3% of participants indicated a higher willingness to outperform their opponents and achieve a higher score when interacting against LLM agents.
In the statistical analysis of Q5, we observed that 83.3% participants expressed more moral condemnation towards the LLM agent than the rule-based AI agent when confronted with betrayal by the opponents. We asked participants who chose the LLM agent about their reasons and summarized several potential underlying factors as follows:
@@
4.4. Semi-structured Interview (RQ3)
To address RQ3, we summarized and organized the content from the interviews, leading to the following conclusions from five distinct perspectives. We revealed the uniqueness of human understanding of LLM interactions through the lens of social cognitive frameworks, as well as the significance of clarifying the LLM agent’s objectives. Additionally, we highlighted participants’ engagement in nurturing the LLM agent, their cautious approach to the agent’s behavior in decision-making scenarios, and the importance of transparency in demonstrating the internal workflows of the LLM agent. These insights underscore the attitudes and behaviors of LLM agents in their interactions with participants.
@@ 4.4.2. Rationalism in Cooperative Behavior Toward Rule-Based AI Agents
When confronted with a purported rule-based AI agent, participants indicated they would begin to betray earlier to probe the opponent’s punishment strategy, thereby adapting their behavior. P4 stated, “It cannot actually obtain money; it is merely following the strategy programmed by its developers.” P14 stated, “I believe I can identify its pattern, and I will formulate my strategy based on its inherent logic.” Participants believed that the rule-based AI agent was predictable and that they could discern the underlying strategy through experimentation. The current perception of rule-based AI in the human mind is that it remains predictable and distinguishable. Consequently, when cooperating with such systems, participants tend to use early variations in strategy to probe the behavior patterns of rule-based AI. This enables them to formulate strategies based on the observed responses, reflecting a behaviorist approach.
@@ 4.4.3. Gradual and Persuasive Guidance: Influencing LLM with My Behavior
4.4.4. Emotion Following Non-cooperative and Its Impact on Subsequent Behavior
Participants attributed the uncooperative behaviors of rule-based AI agents to the deterministic nature of code logic programmed by developers, which explains their indifference towards betrayals by such agents. As P17 aptly remarked, “I don’t feel anything; it’s just the fixed choices of code.” P21 stated, “What algorithm is this giving it decisions? Is it an expert algorithm?.”The content demonstrates that participants exhibit minimal emotional reactions to betrayals by rule-based AI agents, understanding these actions as outcomes of predefined algorithms authored by humans. This underscores a clear distinction in their perceptions and emotional engagements when comparing betrayals by AI agents to those by humans, highlighting the importance of perceived intentionality and human agency in eliciting empathetic and complex emotional responses.
@@ 4.4.4. Emotion Following Non-cooperative and Its Impact on Subsequent Behavior
4.4.5. Perceptions of Different AI Technologies Based on Past Experiences
We examined participants’ experiences and attitudes toward rule-based AI and LLM-based systems across diverse professional fields, including art, mathematics, media, and computer science. Participants previously relied on rule-based, task-specific software for technical tasks, despite its steep learning curves and barriers, exhibiting high trust in its feedback with minimal skepticism. However, the growing adoption of LLM-based technologies has begun to alter this dynamic.
P11 noted, “I feel that nowadays, everything is being labeled as ‘AI-powered,’ and I’m increasingly unsure about what AI is actually doing. This growing prevalence makes me question its reliability.” This characteristic is particularly significant as LLMs substantially reduce the cognitive and temporal costs associated with acquiring new knowledge and skills (Frank, 2023). By leveraging the capabilities of LLMs, users can navigate and assimilate information much more efficiently, thereby optimizing their learning processes. Notably, LLMs have proven to be exceptionally beneficial in aiding participants to understand and apply systems or tools constructed using interpretable algorithms. This facilitation is particularly evident among a diverse group of participants (P3, P5, P9, P11, P13, P20, P24, P25, and P30). These individuals highlighted how LLMs have simplified the complexities associated with these systems, making advanced tools more accessible and easier to comprehend.